OpenStudy (anonymous):

Approximate the root for the following equation using Method of False Position correct upto 5 decimal places. xe^x=cos(x)

OpenStudy (anonymous):

\[f(x)=\cos x-xe^x=0\]\(f(0)=1 >0\) and \(f(1)=\cos 1-e<0\) so there is a root on the interval \((0,1)\)

OpenStudy (anonymous):

sorry it reloaded suddenly

OpenStudy (anonymous):

let \(x_0=0\) , \(x_1=1\) and use the false position rec formula\[x_{2}=\frac{x_0f(x_{1})-x_{1}f(x_0)}{f(x_1)-f(x_0)}\]

OpenStudy (anonymous):

\[x_2= 0.31467\]

OpenStudy (anonymous):

thanks i can do this now:)

OpenStudy (anonymous):

:)